OpenStudy (anonymous):

Solve the quadratic equation by completing the square. x2 + 6x + 7 = 0 a:36+-squareroot2 b:3+-squareroot11 c:-3+-sqaueroot2

OpenStudy (anonymous):

(x+3)^2-2 = 0 Can you do the rest?

OpenStudy (anonymous):

1. Move the number to the opposite side of the equal sign. x2+6x=-7 2. The term with the x, divide in half, then square it. 6/2=3 3^2=9 3. Add that number on both sides of the equation. x2+6x+9=-7+9 4. Solve the perfect square (x+3)^2=2 5. Take the square root on both sides of the equation x+3=+-sqrt(2) 6. Equal the term on the left to the positive and negative roots. Since it's only one term, it's only a matter of solving for x x=-3+-sqrt(2)